# A Theory of Semantic Communication

### v3

@@ VI-A3 Expression and interpretation
$\mathcal{S}_{\text{null}}=\big{\{}\mathcal{UUU},\mathcal{RRR},\mathcal{UUUU},\mathcal{UUUR},\mathcal{UURU},\mathcal{URUU},$
$\mathcal{URRR},\mathcal{RUUU},\mathcal{RURR},\mathcal{RRUR},\mathcal{RRRU},\mathcal{RRRR}\big{\}}$
@@
$\mathcal{S}=\mathcal{S}_{\text{all}}\backslash\mathcal{S}_{\text{null}}=\big{\{}\emptyset,\mathcal{U},\mathcal{R},\mathcal{UU},\mathcal{UR},\mathcal{RU},\mathcal{RR},\mathcal{UUR},\mathcal{URU},$
$\mathcal{RUU},\mathcal{URR},\mathcal{RUR},\mathcal{RRU},\mathcal{UURR}\big{\}},$
@@
$\emptyset$ $1/3$ $2/3$ $1/9$ $1/19$ 00
$\mathcal{U}$ $2/5$ $3/5$ $1/9$ $1/19$ $1$
$\mathcal{R}$ $1/4$ $3/4$ $1/9$ $1/19$ $2$
$\mathcal{UU}$ $1/2$ $1/2$ $1/9$ $1/19$ $2$
$\mathcal{UR}$ $1/3$ $2/3$ $1/9$ $1/19$ $3$
$\mathcal{RU}$ $1/3$ $2/3$ $1/9$ $1/19$ $3$
$\mathcal{RR}$ 00 $1$ 00 $1/19$ $4$
$\mathcal{UUR}$ $1/2$ $1/2$ $1/9$ $1/19$ $4$
$\mathcal{URU}$ $1/2$ $1/2$ $1/9$ $1/19$ $4$
$\mathcal{RUU}$ $1/2$ $1/2$ $1/9$ $1/19$ $4$
$\mathcal{URR}$ 00 $1$ 00 $1/19$ $5$
$\mathcal{RUR}$ 00 $1$ 00 $1/19$ $5$
$\mathcal{RRU}$ 00 $1$ 00 $1/19$ $5$
$\mathcal{UURR}$ 00 $1$ 00 $6/19$ $6$
@@ VI-A4 Semantic encoding
$\mathcal{URR}$ $1$ 00 00 $1$
$\mathcal{RUR}$ $1$ 00 00 $1$
$\mathcal{RRU}$ $1$ 00 00 $1$
$\mathcal{UURR}$ $1$ 00 00 $1$
@@
$\underline{\mathcal{S}}^{\prime}(A)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \underline{\mathcal{S}}^{\prime\prime}(A)=\big{\{}\mathcal{UUR},\mathcal{UURR}\big{\}},$
$\underline{\mathcal{S}}(A)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU},\mathcal{UUR},\mathcal{UURR}\big{\}},$
$\overline{\mathcal{S}}^{\prime}(A)=\big{\{}\emptyset,\mathcal{RR}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\mathcal{S}}^{\prime\prime}(A)=\big{\{}\mathcal{UURR}\big{\}},$
$\overline{\mathcal{S}}(A)=\big{\{}\emptyset,\mathcal{RR},\mathcal{UURR}\big{\}},$
$\underline{\mathcal{S}}^{\prime}(B)=\big{\{}\emptyset,\mathcal{RR}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \underline{\mathcal{S}}^{\prime\prime}(B)=\big{\{}\mathcal{UURR}\big{\}},$
$\underline{\mathcal{S}}(B)=\big{\{}\emptyset,\mathcal{RR},\mathcal{UURR}\big{\}},$
$\overline{\mathcal{S}}^{\prime}(B)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\mathcal{S}}^{\prime\prime}(B)=\big{\{}\mathcal{UUR},\mathcal{UURR}\big{\}},$
$\overline{\mathcal{S}}(B)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU},\mathcal{UUR},\mathcal{UURR}\big{\}},$
@@
$D_{\underline{\bm{U}}^{(0)},\bm{Q}}$ $=$ $p(A)\varphi(A,\emptyset)+p(B)\varphi(B,\emptyset)=\frac{4}{9},$
$L_{\underline{\bm{U}}^{(0)}}$ $=$ $p(A)\ell(\emptyset)+p(B)\ell(\emptyset)=0.$
@@
$G(w=A,\emptyset,\mathcal{U})=-\frac{1}{15},$
$G(w=A,\emptyset,\mathcal{UU})=-\frac{1}{12},$
$G(w=B,\emptyset,\mathcal{RR})=-\frac{1}{12}.$
@@
$D_{\underline{\bm{U}}^{(1)},\bm{Q}}$ $=$ $p(A)\varphi(A,\mathcal{UU})+p(B)\varphi(B,\emptyset)=\frac{7}{18},$
$L_{\underline{\bm{U}}^{(1)}}$ $=$ $p(A)\ell(\mathcal{UU})+p(B)\ell(\emptyset)=\frac{2}{3}.$
@@
      $\underline{U}^{(0)}=\big{\{}\emptyset,\emptyset\big{\}}$       $4/9$       00
@@
      $\underline{U}^{(3)}=\big{\{}\mathcal{UUR},\mathcal{RR}\big{\}}$       $1/6$       $4$
@@
      $\underline{U}^{(5)}=\big{\{}\mathcal{UURR},\mathcal{UURR}\big{\}}$       $1/3$       $6$
      $\overline{U}^{(0)}=\big{\{}\emptyset,\emptyset\big{\}}$       $4/9$       00
@@
      $\overline{U}^{(5)}=\big{\{}\mathcal{UURR},\mathcal{UURR}\big{\}}$       $1/3$       $6$
@@ VI-A5 Semantic decoding and CSED
Under the Hamming distortion, we have $\psi_{p}(\hat{w},s)=p(s)-p(\hat{w})p(s|\hat{w})$. The semantic decoding schemes $\widetilde{\bm{\Delta}}_{n^{\prime}_{1},n^{\prime}_{2},...,n^{\prime}_{M}}$ and $\widetilde{\bm{\Delta}}_{n^{\prime\prime}_{1},n^{\prime\prime}_{2},...,n^{\prime\prime}_{M}}$ can be derived from Theorem 5.1, as listed in Table IV, and
$D_{\widetilde{\bm{\Delta}}_{n^{\prime}_{1},n^{\prime}_{2},...,n^{\prime}_{M}}}\approx 0.3158,$ $D_{\widetilde{\bm{\Delta}}_{n^{\prime\prime}_{1},n^{\prime\prime}_{2},...,n^{\prime\prime}_{M}}}\approx 0.6842,$
$L_{\bm{P}}\approx 3.5887.$
@@ VII-B Multiple Transmissions
    Transmission of infinite meanings. With an infinite number of meanings to be transmitted, multiple transmissions provide a significant advantage by allowing the receiver to gather $p(\hat{s})$ for the received symbols, thereby facilitating the estimation of $p(s)$ and $p(w)$ via either direct inversion or regularization techniques, such as LASSO and Tikhonov regularization. In this scenario: i) Following the semantic decoding approach outlined in Proposition 5.2, we can achieve the optimal distortion $D_{\bm{P},\bm{V}^{*}_{p}}$, optimizing the fidelity of communicated meanings. ii) The receiver can leverage the concept of typicality in the transmitted sequence [29], focusing on decoding among sequences that are “typical”, given the observed transmissions. This approach takes advantage of statistical regularities in the sequence of transmitted meanings, enhancing decoding efficiency and accuracy.
@@ VIII Conclusion
$D_{\bm{U},\bm{Q}}=\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}u(s_{m}|w_{n})q(\hat{w}|s_{m})$
$=\sum_{i_{1},i_{2},...,i_{N}}u(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})$
×$\times\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}\delta_{i_{1},i_{2},...,i_{N}}(s_{m}|w_{n})q(\hat{w}|s_{m})$
$=\sum_{i_{1},i_{2},...,i_{N}}u(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})D_{{\bm{\Delta}}_{i_{1},i_{2},...,i_{N}},\bm{Q}}.$
@@
$L_{\bm{U}}=\sum_{n}p(w_{n})\sum_{m}u(s_{m}|w_{n})\ell(s_{m})$
$=\sum_{i_{1},i_{2},...,i_{N}}u(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})$
×$\times\sum_{n}p(w_{n})\sum_{m}\delta_{i_{1},i_{2},...,i_{N}}(s_{m}|w_{n})\ell(s_{m})$
$=\sum_{i_{1},i_{2},...,i_{N}}u(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})L_{{\bm{\Delta}}_{i_{1},i_{2},...,i_{N}}}.$
@@
$\sum_{s}u(s|w)$ $=$ $\sum_{s}\sum_{i=1}^{|\bm{A}|}\lambda_{i}{\bm{\Delta}}_{i}(s|w)=\sum_{i=1}^{|\bm{A}|}\lambda_{i}\sum_{s}{\bm{\Delta}}_{i}(s|w)$
$=$ $\sum_{i=1}^{|\bm{A}|}\lambda_{i}\times 1=1.$
@@
$\begin{split}L_{\bm{U}}&=\sum_{n}p(w_{n})\sum_{m}u(s_{m}|w_{n})\ell(s_{m})\\ &=\sum_{n}p(w_{n})\sum_{m}\sum_{i=1}^{|\bm{A}|}\lambda_{i}{\bm{\Delta}}_{i}(s|w)\ell(s_{m})\\ &=\sum_{i=1}^{|\bm{A}|}\lambda_{i}\sum_{n}p(w_{n})\sum_{m}{\bm{\Delta}}_{i}(s|w)\ell(s_{m})\\ &=\sum_{i=1}^{|\bm{A}|}\lambda_{i}L_{{\bm{\Delta}}_{i}}=L\\ \end{split}$
@@
$D_{\bm{U},\bm{Q}}=\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}u(s_{m}|w_{n})q(\hat{w}|s_{m})$
$=\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}\sum_{i=1}^{|\bm{A}|}\lambda_{i}{\bm{\Delta}}_{i}(s|w)q(\hat{w}|s_{m})$
$=\sum_{i=1}^{|\bm{A}|}\lambda_{i}\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}{\bm{\Delta}}_{i}(s|w)q(\hat{w}|s_{m})$
$=\sum_{i=1}^{|\bm{A}|}\lambda_{i}D_{{\bm{\Delta}}_{i},\bm{Q}}=D.$
@@ Appendix B Properties of the Six Subsets
${\underline{m}}_{w,1}=\min{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime}\right\}},{\underline{m}}_{w,2}=\max{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime}\right\}},$
${\underline{m}}_{w,3}=\min{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime\prime}\right\}},{\underline{m}}_{w,4}=\max{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime\prime}\right\}},$
${\overline{m}}_{w,1}=\min{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime}\right\}},{\overline{m}}_{w,2}=\max{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime}\right\}},$
${\overline{m}}_{w,3}=\min{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime\prime}\right\}},{\overline{m}}_{w,4}=\max{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime\prime}\right\}}.$
@@
For any $n=1,2,...,N$, $t=1,2,...,\underline{T}$ and $s_{m}\in{\underline{\mathcal{S}}}_{w_{n}}$,
@@
${\underline{G}}^{(\tilde{t})}=G\left(w_{n},s_{{\underline{i}}_{n}^{(\tilde{t}-1)}},s_{{\underline{i}}_{n}^{(\tilde{t})}}\right).$
@@
${\underline{G}}^{(\tilde{t})}\geq G\left(w_{n},s_{m},s_{{\underline{i}}_{n}^{(\tilde{t})}}\right).$ (51)
@@
${\underline{t}}^{\prime}\triangleq\min\left\{t:L_{{\underline{{\bm{U}}}}^{(t)}}\geq L^{\prime}\right\},{\underline{t}}^{\prime\prime}\triangleq\min\left\{t:\allowbreak L_{{\underline{{\bm{U}}}}^{(t)}}\allowbreak\geq L^{\prime\prime}\right\}$
$\overline{t}^{\prime}\triangleq\min\left\{t:L_{{\overline{{\bm{U}}}}^{\left({\overline{t}}^{\prime}\right)}}\geq L^{\prime}\right\},\overline{t}^{\prime\prime}\triangleq\min\left\{t:L_{{\overline{{\bm{U}}}}^{(t)}}\geq L^{\prime\prime}\right\}.$
@@
$\underline{D}(L^{\prime\prime})=\underline{D}(L^{\prime})+\left(\underline{D}_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-\underline{D}(L^{\prime})\right)$
$+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}\left({\underline{D}}_{{\underline{{\bm{U}}}}^{(t)}}-{\underline{D}}_{{\bm{U}}^{\left(t-1\right)}}\right)+\left(\underline{D}(L^{\prime\prime})-{\underline{D}}_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}-1\right)}}\right)$
$=\underline{D}(L^{\prime})+{\underline{J}}^{({\underline{t}}^{\prime})}\left(L_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-L^{\prime}\right)$
$+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}{{\underline{J}}^{(t)}\left(L_{{\underline{{\bm{U}}}}^{(t)}}-L_{{\underline{{\bm{U}}}}^{\left(t-1\right)}}\right)}+{\underline{J}}^{\left(\underline{t}^{\prime\prime}\right)}\left(L^{\prime\prime}-L_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}-1\right)}}\right).$

### v4

@@ VI-A3 Expression and interpretation
$\hskip-14.22636pt\mathcal{S}_{\text{null}}=\big{\{}\mathcal{UUU},\mathcal{RRR},\mathcal{UUUU},\mathcal{UUUR},\mathcal{UURU},\mathcal{URUU},$
$\hskip-14.22636pt\mathcal{URRR},\mathcal{RUUU},\mathcal{RURR},\mathcal{RRUR},\mathcal{RRRU},\mathcal{RRRR}\big{\}}$
@@
$\hskip-14.22636pt\mathcal{S}=\mathcal{S}_{\text{all}}\backslash\mathcal{S}_{\text{null}}=\big{\{}\emptyset,\mathcal{U},\mathcal{R},\mathcal{UU},\mathcal{UR},\mathcal{RU},\mathcal{RR},\mathcal{UUR},\mathcal{URU},$
$\hskip-14.22636pt\mathcal{RUU},\mathcal{URR},\mathcal{RUR},\mathcal{RRU},\mathcal{UURR}\big{\}},$
@@
$\emptyset$ $1/3$ $2/3$ $1/9$ $1/19$ 0
$\mathcal{U}$ $2/5$ $3/5$ $1/9$ $1/19$ 11
$\mathcal{R}$ $1/4$ $3/4$ $1/9$ $1/19$ 22
$\mathcal{UU}$ $1/2$ $1/2$ $1/9$ $1/19$ 22
$\mathcal{UR}$ $1/3$ $2/3$ $1/9$ $1/19$ 33
$\mathcal{RU}$ $1/3$ $2/3$ $1/9$ $1/19$ 33
$\mathcal{RR}$ 0 11 0 $1/19$ 44
$\mathcal{UUR}$ $1/2$ $1/2$ $1/9$ $1/19$ 44
$\mathcal{URU}$ $1/2$ $1/2$ $1/9$ $1/19$ 44
$\mathcal{RUU}$ $1/2$ $1/2$ $1/9$ $1/19$ 44
$\mathcal{URR}$ 0 11 0 $1/19$ 55
$\mathcal{RUR}$ 0 11 0 $1/19$ 55
$\mathcal{RRU}$ 0 11 0 $1/19$ 55
$\mathcal{UURR}$ 0 11 0 $6/19$ 66
@@ VI-A4 Semantic encoding
$\mathcal{URR}$ 11 0 0 11
$\mathcal{RUR}$ 11 0 0 11
$\mathcal{RRU}$ 11 0 0 11
$\mathcal{UURR}$ 11 0 0 11
@@
$\hskip-14.22636pt\underline{\mathcal{S}}^{\prime}(A)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \underline{\mathcal{S}}^{\prime\prime}(A)=\big{\{}\mathcal{UUR},\mathcal{UURR}\big{\}},$
$\hskip-14.22636pt\underline{\mathcal{S}}(A)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU},\mathcal{UUR},\mathcal{UURR}\big{\}},$
$\hskip-14.22636pt\overline{\mathcal{S}}^{\prime}(A)=\big{\{}\emptyset,\mathcal{RR}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\mathcal{S}}^{\prime\prime}(A)=\big{\{}\mathcal{UURR}\big{\}},$
$\hskip-14.22636pt\overline{\mathcal{S}}(A)=\big{\{}\emptyset,\mathcal{RR},\mathcal{UURR}\big{\}},$
$\hskip-14.22636pt\underline{\mathcal{S}}^{\prime}(B)=\big{\{}\emptyset,\mathcal{RR}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \underline{\mathcal{S}}^{\prime\prime}(B)=\big{\{}\mathcal{UURR}\big{\}},$
$\hskip-14.22636pt\underline{\mathcal{S}}(B)=\big{\{}\emptyset,\mathcal{RR},\mathcal{UURR}\big{\}},$
$\hskip-14.22636pt\overline{\mathcal{S}}^{\prime}(B)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU}\big{\}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\mathcal{S}}^{\prime\prime}(B)=\big{\{}\mathcal{UUR},\mathcal{UURR}\big{\}},$
$\hskip-14.22636pt\overline{\mathcal{S}}(B)=\big{\{}\emptyset,\mathcal{U},\mathcal{UU},\mathcal{UUR},\mathcal{UURR}\big{\}},$
@@
$D_{\underline{\bm{U}}^{(0)},\bm{Q}}\hskip-5.69046pt$ $=$ $\hskip-5.69046ptp(A)\varphi(A,\emptyset)+p(B)\varphi(B,\emptyset)=\frac{4}{9},$
$L_{\underline{\bm{U}}^{(0)}}\hskip-5.69046pt$ $=$ $\hskip-5.69046ptp(A)\ell(\emptyset)+p(B)\ell(\emptyset)=0.$
@@
$\hskip-14.22636ptG(w=A,\emptyset,\mathcal{U})=-\frac{1}{15},$
$\hskip-14.22636ptG(w=A,\emptyset,\mathcal{UU})=-\frac{1}{12},$
$\hskip-14.22636ptG(w=B,\emptyset,\mathcal{RR})=-\frac{1}{12}.$
@@
$D_{\underline{\bm{U}}^{(1)},\bm{Q}}\hskip-5.69046pt$ $=$ $\hskip-5.69046ptp(A)\varphi(A,\mathcal{UU})+p(B)\varphi(B,\emptyset)=\frac{7}{18},$
$L_{\underline{\bm{U}}^{(1)}}\hskip-5.69046pt$ $=$ $\hskip-5.69046ptp(A)\ell(\mathcal{UU})+p(B)\ell(\emptyset)=\frac{2}{3}.$
@@
  $\underline{U}^{(0)}=\big{\{}\emptyset,\emptyset\big{\}}$   $4/9$   0
@@
  $\underline{U}^{(3)}=\big{\{}\mathcal{UUR},\mathcal{RR}\big{\}}$   $1/6$   44
@@
  $\underline{U}^{(5)}=\big{\{}\mathcal{UURR},\mathcal{UURR}\big{\}}$   $1/3$   66
  $\overline{U}^{(0)}=\big{\{}\emptyset,\emptyset\big{\}}$   $4/9$   0
@@
  $\overline{U}^{(5)}=\big{\{}\mathcal{UURR},\mathcal{UURR}\big{\}}$   $1/3$   66
@@ VI-A5 Semantic decoding and CSED
[…]
$D_{\widetilde{\bm{\Delta}}_{n^{\prime}_{1},n^{\prime}_{2},...,n^{\prime}_{M}}}\approx 0.3158,$ $\hskip-11.38092ptD_{\widetilde{\bm{\Delta}}_{n^{\prime\prime}_{1},n^{\prime\prime}_{2},...,n^{\prime\prime}_{M}}}\approx 0.6842,$
$\hskip-42.67912ptL_{\bm{P}}\approx 3.5887.$
@@ VII-B Multiple Transmissions
    Transmission of infinite meanings. With an infinite number of meanings to be transmitted, multiple transmissions provide a significant advantage by allowing the receiver to gather p(s^)p($\hat${s}) for the received symbols, thereby facilitating the estimation of $p(s)$ and $p(w)$ via either direct inversion or regularization techniques, such as LASSO and Tikhonov regularization. In this scenario: i) Following the semantic decoding approach outlined in Proposition 5.2, we can achieve the optimal distortion $D_{\bm{P},\bm{V}^{*}_{p}}$, optimizing the fidelity of communicated meanings. ii) The receiver can leverage the concept of typicality in the transmitted sequence [29], focusing on decoding among sequences that are “typical”, given the observed transmissions. This approach takes advantage of statistical regularities in the sequence of transmitted meanings, enhancing decoding efficiency and accuracy.
@@ VIII Conclusion
$\hskip-18.49411ptD_{\bm{U},\bm{Q}}=\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}u(s_{m}|w_{n})q(\hat{w}|s_{m})$
$\hskip-18.49411pt=\sum_{i_{1},i_{2},...,i_{N}}u(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})$
$\hskip-18.49411pt\times\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}\delta_{i_{1},i_{2},...,i_{N}}(s_{m}|w_{n})q(\hat{w}|s_{m})$
$\hskip-18.49411pt=\hskip-5.69046pt\sum_{i_{1},i_{2},...,i_{N}}\hskip-5.69046ptu(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})D_{{\bm{\Delta}}_{i_{1},i_{2},...,i_{N}},\bm{Q}}.$
@@
$\hskip-18.49411ptL_{\bm{U}}=\sum_{n}p(w_{n})\sum_{m}u(s_{m}|w_{n})\ell(s_{m})$
$\hskip-18.49411pt=\sum_{i_{1},i_{2},...,i_{N}}u(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})$
$\hskip-18.49411pt\times\sum_{n}p(w_{n})\sum_{m}\delta_{i_{1},i_{2},...,i_{N}}(s_{m}|w_{n})\ell(s_{m})$
$\hskip-18.49411pt=\sum_{i_{1},i_{2},...,i_{N}}\hskip-5.69046ptu(s_{i_{1}}|w_{1})u(s_{i_{2}}|w_{2})...u(s_{i_{N}}|w_{N})L_{{\bm{\Delta}}_{i_{1},i_{2},...,i_{N}}}.$
@@
$\sum_{s}u(s|w)\hskip-5.69046pt$ $=$ $\hskip-5.69046pt\sum_{s}\sum_{i=1}^{|\bm{A}|}\lambda_{i}{\bm{\Delta}}_{i}(s|w)=\sum_{i=1}^{|\bm{A}|}\lambda_{i}\sum_{s}{\bm{\Delta}}_{i}(s|w)$
$=$ $\hskip-5.69046pt\sum_{i=1}^{|\bm{A}|}\lambda_{i}\times 1=1.$
@@
[…]
@@
$\hskip-18.49411ptD_{\bm{U},\bm{Q}}=\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}u(s_{m}|w_{n})q(\hat{w}|s_{m})$
$\hskip-18.49411pt=\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}\sum_{i=1}^{|\bm{A}|}\lambda_{i}{\bm{\Delta}}_{i}(s|w)q(\hat{w}|s_{m})$
$\hskip-18.49411pt=\sum_{i=1}^{|\bm{A}|}\lambda_{i}\sum_{n}p(w_{n})\sum_{\hat{w}}d(w_{n},\hat{w})\sum_{m}{\bm{\Delta}}_{i}(s|w)q(\hat{w}|s_{m})$
$\hskip-18.49411pt=\sum_{i=1}^{|\bm{A}|}\lambda_{i}D_{{\bm{\Delta}}_{i},\bm{Q}}=D.$
@@ Appendix B Properties of the Six Subsets
$\hskip-18.49411pt{\underline{m}}_{w,1}=\min{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime}\right\}},{\underline{m}}_{w,2}=\max{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime}\right\}},$
$\hskip-18.49411pt{\underline{m}}_{w,3}=\min{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime\prime}\right\}},{\underline{m}}_{w,4}=\max{\left\{m:s_{m}\in{\underline{\mathcal{S}}}_{w}^{\prime\prime}\right\}},$
$\hskip-18.49411pt{\overline{m}}_{w,1}=\min{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime}\right\}},{\overline{m}}_{w,2}=\max{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime}\right\}},$
$\hskip-18.49411pt{\overline{m}}_{w,3}=\min{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime\prime}\right\}},{\overline{m}}_{w,4}=\max{\left\{m:s_{m}\in{\overline{\mathcal{S}}}_{w}^{\prime\prime}\right\}}.$
@@
For any n=1,2,,Nn=1,2,...,N, $t=1,2,...,\underline{T}$ and $s_{m}\in{\underline{\mathcal{S}}}_{w_{n}}$,
@@
[…]
@@
[…]
@@
$\hskip-18.49411pt{\underline{t}}^{\prime}\triangleq\min\left\{t:L_{{\underline{{\bm{U}}}}^{(t)}}\geq L^{\prime}\right\},{\underline{t}}^{\prime\prime}\triangleq\min\left\{t:\allowbreak L_{{\underline{{\bm{U}}}}^{(t)}}\allowbreak\geq L^{\prime\prime}\right\}$
$\hskip-18.49411pt\overline{t}^{\prime}\triangleq\min\left\{t:L_{{\overline{{\bm{U}}}}^{\left({\overline{t}}^{\prime}\right)}}\geq L^{\prime}\right\},\overline{t}^{\prime\prime}\triangleq\min\left\{t:L_{{\overline{{\bm{U}}}}^{(t)}}\geq L^{\prime\prime}\right\}.$
@@
$\hskip-18.49411pt\underline{D}(L^{\prime\prime})=\underline{D}(L^{\prime})+\left(\underline{D}_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-\underline{D}(L^{\prime})\right)$
$\hskip-18.49411pt+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}\left({\underline{D}}_{{\underline{{\bm{U}}}}^{(t)}}-{\underline{D}}_{{\bm{U}}^{\left(t-1\right)}}\right)+\left(\underline{D}(L^{\prime\prime})-{\underline{D}}_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}-1\right)}}\right)$
$\hskip-18.49411pt=\underline{D}(L^{\prime})+{\underline{J}}^{({\underline{t}}^{\prime})}\left(L_{{\underline{{\bm{U}}}}^{(\underline{t}^{\prime})}}-L^{\prime}\right)$
$\hskip-18.49411pt+\sum_{t=\underline{t}^{\prime}+1}^{\underline{t}^{\prime\prime}-1}{{\underline{J}}^{(t)}\left(L_{{\underline{{\bm{U}}}}^{(t)}}-L_{{\underline{{\bm{U}}}}^{\left(t-1\right)}}\right)}+{\underline{J}}^{\left(\underline{t}^{\prime\prime}\right)}\left(L^{\prime\prime}-L_{{\underline{{\bm{U}}}}^{\left(\underline{t}^{\prime\prime}-1\right)}}\right).$